OpenStudy (anonymous):

Solve the equation. 4x^2 - 20 = 5

OpenStudy (anonymous):

4x² - 20 = 5 4x² - 25 = 0 Now factor 4x² - 25.

OpenStudy (anonymous):

\[4x ^{2}-20=5\]\[4x ^{2}-20+20=5+25\]\[4x ^{2}=25\]\[x ^{2}=\frac{ 25 }{ 4 }\]\[x=\sqrt{\frac{ 25 }{ 4 }}\]\[x=\frac{ 5 }{ 2 }\]\[x=2\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

@EchaSetiaWijaya X is also -2.5. Remember, √(x²) = |x|

OpenStudy (anonymous):

yes, thanks for correction